Ricks Paint Store sells a special shade of pink paint (Ricky's Blush) which can be made from red paint and white paint. The amount of red paint should be three times the amount of white paint. How much of each kind of paint is needed to make 264 gallons of Ricky's Blush?
Please help me

Answer by nerdybill(7384)

Ricks Paint Store sells a special shade of pink paint (Ricky's Blush) which can be made from red paint and white paint. The amount of red paint should be three times the amount of white paint. How much of each kind of paint is needed to make 264 gallons of Ricky's Blush?
.
Let x = amount (gals) of white paint
then
3x = amount (gals) of red paint
.
x + 3x = 264
4x = 264
x = 264/4
x = 66 gals (white paint)
.
Red paint:
3x = 3(66) = 198 gals